OpenStudy (anonymous):

Let u = <3, -1>, v = <-6, -6>. Find 9u + 2v.

OpenStudy (jdoe0001):

\(\large \begin{cases} 9\cdot <a,b>\to <9\cdot a,9\cdot b>\\ 2\cdot <a,b>\to <2\cdot a,2\cdot b>\\ \hline\\ <a,b>+<c,d>\to <a+c, b+d> \end{cases}\)

OpenStudy (anonymous):

<27, -9> +<-12,-12> <15,-21> Is that right?

OpenStudy (jdoe0001):

<15,-21> \(\huge \checkmark\)